OpenStudy (anonymous):

y= 2-x and y= 4x+7 Answers: A: 1,1 B: 1,3 C: -1,3 D: -1,1 E: no solution

OpenStudy (texaschic101):

y = 2 - x so sub in 2 - x in place of y in the other equation y = 4x + 7 2 - x = 4x + 7 --- add x to both sides 2 = 4x + x + 7 2 = 5x + 7 --- subtract 7 from both sides 2 - 7 = 5x -5 = 5x --- divide both sides by 5 -5/5 = x -1 = x now sub -1 in for x in either equation to find y y = 2 - x y = 2 - (-1) y = 2 + 1 y = 3 check.. y = 4x + 7 3 = 4(-1) + 7 3 = -4 + 7 3 = 3 (correct) so your solution is : (-1,3)